Solve Equation
ln(3x+1)/ln(2x-1)=2

Answer by lwsshak3(6504)

Solve Equation
ln(3x+1)/ln(2x-1)=2
ln(3x+1)=2ln(2x-1)
ln(3x+1)=ln(2x-1)^2
3x+1=(2x-1)^2
3x+1=4x^2-4x+1
4x^2-7x=0
x(4x-7)=0
x=0 (reject,(2x-1)>0)
x=7/4